### v1

# Hidden long-range correlations in the ion distribution at the graphite / [bmim][NTf2] electrified interface

Abstract
A capacitor consisting of the [bmim][NTf2] ionic liquid (IL) confined in between planar graphite electrodes has been investigated by molecular dynamics based on an all-atom, unpolarizable force field. Structural and dynamical properties such as: (i) the density and orientation of the [bmim]+ and [NTf2]- ions throughout the capacitor; (ii) the electrostatic double layer at the electrode / electrolyte interface; (iii) the ions’ mobility perpendicular and parallel to the graphite plates are determined as a function of the electrostatic charge of the capacitor, the concentration of absorbed water, the temperature and pressure. Grouping the [bmim]+ and [NTf2]- ions into neutral ion pairs reveals an intriguing ordering normal to the interface that is related to correlations among the dipole moments of the neutral ion pairs. These correlations might explain the observation of an anomalous Stark effect (Pockels effect) reported a few years ago in Langmuir 2021, 37, 5193-5201, and provides useful insight for the multitude of electro-chemical applications that involve electrode / ionic liquid interfaces.
Corresponding authors: corteshu@mpip-mainz.mpg.de, nforerom@uni-mainz.de
I Introduction
The electrostatic double layer is a unifying concept that plays a fundamental role in the chemical-physics of interfaces. As such, it also represents a crucial aspect of most electro-chemical devices and related processes.bock In the last two decades the interest in this time honored topic has been enhanced by two developments, i.e., the expanded role of electro-chemical energy conversion and storage applications (batteries, fuel cells, super-capacitors) as well as the surge of activity concerning ionic liquids (ILs) which offer tantalizing opportunities for increasing efficiency, durability, safety and environmental compatibility of electrochemical processes and devices. References on those two research fields are too numerous to cite, and we point to the extensive discussions in Ref. korn1, ; korn2, ; korn3, .
Figure 1: The dry graphite/[bmim][NTf2]/graphite capacitor. Gray, blue, red, green and yellow dots represent C, N, O, F and S atoms, respectively. Hydrogen atoms not shown. The regular vertical layers of C atoms are part of the graphite electrodes. The right-most graphite layer in the figure is the periodic replica of the left-most layer.
The long range of the Coulomb charge-charge interactions poses strict conditions to the structure and dynamics of simple electrolytes in contact with an electrode. In particular, the width of the spatial charge separation at the interface is limited by screening to a few characteristic lengths, i.e., multiples of a screening length $\lambda$, whose value depends primarily on the thermodynamic state of the electrolyte. Beyond this (usually microscopic) length, the electrolyte appears to be largely bulk-like.
Given this generally accepted picture, it came as a surprise the revelation by optical measurements of characteristic shifts in the frequency of characteristic lines, affecting the spectroscopic properties of IL electrolytes far from the interface.expLangmuir
Since the strict confinement of the interfacial charge perturbation is supported by very robust arguments and is verified by a large number of experimental, theoretical and computational studies, we look for the responsible of this long range ordering at the second simpler and most likely suspect, represented by the organisation of dipoles by long range correlation effects. As the testbed for this analysis we have chosen the interface between a graphite electrode and a paradigmatic IL, consisting of the 1-butyl-3-methylimidazolium cation ([bmim]+) and bis(trifluoromethylsulfonyl)-imide anion ([NTf2]-), i.e., [bmim][NTf2]. This compound is mildly hydrophobic but retains a not-completely-negligible water concentration of up to about $25$ molar %, which however corresponds to only less that $1.5$ wgt% due to the mass disparity of water and [bmim][NTf2]. We chose this system for our investigation, despite its complexity, because of its interest as a solvent electrolyteelectrochem and relatively abundance of experimental data on its physical chemistry properties. Moreover, the same effect might not be observable in simpler (either real or idealised) models because size and shape of the electrolyte ions might play a role in the propagation of correlation effects far from the interface.
The first target of this investigation is to define a population of instantaneous dipoles made by the IL ions in the least arbitrary way. Achieving this goal will allow us to quantify correlations among dipoles as a function of the electron density on the electrode, of temperature and of concentration of water contamination, due to the hygroscopic character of [bmim][NTf2]. The MD of the interface, that represents the required background of this investigation, provides a wealth of additional information on the electrostatic properties of the interface, including an estimation of the differential capacitance of the interface, and its dependence on temperature, pressure, surface charge density on the electrode, and concentration of contaminant water. Since the discussion of all these aspects could obscure the major novelty of the present study, we defer the full presentation of our investigation to a later and more comprehensive paper, devoting the present document to a concise report of the main result, represented by the presence of dipole correlations far from the interface, where the average charge density and electric field vanish.
II Model and Method
The simulated system is a planar capacitor consisting of two parallel graphite plates orthogonal to the $\hat{x}$ axis, delimiting a relatively thick ($\sim 11$ nm) [bmim][NTf2] layer, see Fig. 1. Both graphite and the IL are described at the atomistic level. The potential energy surface of the system is represented through an empirical force field of the Amber form,amber parametrised at first using the Gromos model in its 54a754a7 version for the [bmim][NTf2] component, gromos and the SPC model for water.spc In practice, the topology files for the [bmim]+ and [NTf2]- ions have been obtained using the automated topology building utility made available by the ATB website.atb At first, all the charges on the graphite carbon atoms are set to zero.
The force field for graphite has been built by generating the topology files for coronene (C24H12) using again Gromos 54a7 and ATB, then copying the stretching, bending, dihedral and improper dihedral parameters for aromatic carbon to the graphite topology file. The 2D graphite periodicity along the plane of the electrode is built into the topology file by listing all the bonded C pairs including those that connect through pbc.
Figure 2: (a) Honeycomb structure of the single graphite layer (graphene). (b) stacking of A (red) and B (black) layers whose repetition is makes the finite graphite slab representing the planar plates of the capacitor. Slight irregularities in the atomic positions are due to the fact that the configuration has been taken from an MD simulation at $T=300$ K.
Test simulations of homogeneous (3D periodic boundary conditions, pbc) [bmim][NTf2] systems have shown that the self-diffusion constant of both [bmim]+ and [NTf2]- as computed by the Einstein relation are severely underestimated. This is a know problem of unpolarisable IL force fields,diffu such as Gromos 54a7, and, in the present case the problem might be exacerbated by a non-negligible over-estimation of the homogeneous IL density. The diffusivity problem sometimes is solved by scaling all atomic charges by a 0.80.90.8-0.9 factor,08 ; 09 justified in terms of atomic polarizability or partial charge transfer among cations and anions (which can also be seen as a form of electronic polarizability). These effects, however, are likely to be different in the bulk and at an interface. Moreover, any ion that (hypothetically) leaves the condensed phase needs to have an integer charge. Lacking a suitable model to include the charge dependence on configuration and thermodynamic state in a consistent and computationally convenient form, it has been decided to keep unchanged the formal charge +1e+1e, 1e-1e of cations and anions, respectively. This decision implies accepting a general slowing down of the ion dynamics with respect to the experimental system. The density over-estimation, instead, has been addressed by an empirical tuning of the Lennard-Jones (LJ) pair potential, hoping to improve at the same time the diffusivity aspect. To this aim, all Lennard-Jones radii $\sigma_{ij}$ of [bmim][NTf2] atoms have been expanded by a constant factor of $1.03$ (+3%), recovering the experimental density of pure and homogeneous [bmim][NTf2] at $T=300$ K. The homogeneous liquid properties of [bmim][NTf2] estimated by this model are reported and briefly discussed in Sec. III. The aromatic carbon description of graphite has been left unchanged at the Gromos 54a7 level, since it has little effect on the IL density and dynamics even in its proximity besides confining the IL to the inside of the model capacitor.
Figure 3: [bmim][NTf2] neutral ion pair. The picture shows a generic configuration from a MD trajectory at $T=300$ K.
[…]
All planar capacitor samples are enclosed in an orthorhombic box, with periodic boundary conditions applied. More precisely, at first samples are periodically replicated in 2D along the $(yz)$ plane. Then, despite the inherent 2D character of the capacitor, the electrostatic interactions have been computed using 3D Ewald sums upon replicating the capacitor also along the $\hat{x}$ direction, because 2D Ewald sum routines are generally considered to be less efficient than their 3D counterparts. No explicit test, however, has been carried out to quantify or even verify this statement for the present case. The thickness of the graphite slab ($\sim 30$ Å ) limits the spurious interactions of the two interfaces that, because of pbc along $x$ find themselves weakly interacting through each graphite block.
The artificial periodicity along $x$, however, has a far more significant effect, that needs to be accounted for. An isolated, single planar capacitor of charge $Q=\sigma A$, where $\sigma$ is the surface charge density on the positive electrode and $A$ is the electrode area, is globally neutral, and the electric field on its outside vanishes. Hence, application of Gauss’ theorem immediately gives that the electric field strength on its inside is a constant equal to $4\pi\sigma$. For an infinitely periodic sequence of capacitors, however, the analysis is not so simple, because there is no simple way to locate a surface of vanishing electric field (${\bf E}=0$) from which to start applying Gauss’ theorem to compute ${\bf E}(x)$ anywhere else in space. For this reason, the electric field inside the capacitor, although still constant and directed along $x$, is not equal to $4\pi\sigma$.chris The precise value of ${\bf E}(x)$ can be computed analytically based on the Ewald sum formalism, but it depends on boundary condition parameters set in the simulation code, whose value in not easily determined. To avoid this ambiguity, the electric field $^{\bf E_{0}}=E_{0}(x)\hat{x}$ along the $x$ coordinate within the capacitor and within the graphite slab has been computed using Gromacs considering an empty capacitor whose charged plates (represented by the charge graphite layers) are located at the distance used in the actual simulations with the ionic liquid. This empty capacitor curve $E_{0}(x)$ has been used throughout the paper to compute the position-dependent electric field ${\bf E}(x)$ and electrostatic potential ${\bf\phi}(x)$ through the capacitor.
Most simulations have been carried out in the NPT ensemble, using the Parrinello-Rahman approach. The remaining few simulations have been done in the NVT ensemble, mainly to have a fixed reference frame to compute and plot density distribution functions perpendicular to the interface. The NVT ensemble has also been used to simulate free standing dry and wet ionic liquid slabs, exposing two parallel (on average) IL/vacuum interfaces.
The graphite plates, in particular, consist of hexagonal layers of carbon, with aromatic sp2 bonds among them. A solid-like graphite slab is formed considering a stack of 9 layers in the ABABABABA configuration (see Fig. 2). It has been verified that the NPT simulation of the identically neutral 9-layers graphite slab in pbc at 1 bar (no vacuum) and $T=300$ K reproduces fairly well the structural parameters of graphite, included the wide layer-layer separation $d\sim 3.5$ Å .
All simulated systems consist of the same number of cations and anions. The electrolyte slab, therefore, is globally neutral. Since also the whole sample, which is 3D periodic, needs to be neutral, the two interfaces delimiting the electrolyte may only carry zero or opposite surface charge densities $\pm\sigma$. At first, the charge on each graphite C atom is identically zero. Then, any non-vanishing charge density on the graphite surfaces will be represented by atomic charges $q_{grf}$ of equal modulus and opposite sign on the two surfaces. Since graphite is a conductor, the static $\pm\sigma$ charges on the two sides of the graphite, in reality, would not be stable with respect to charge recombination. However, the electrical conductivity perpendicular to the graphite planes is about $10^{3}$ times lower than along the planes. Therefore, the simple charging model is justified by neglecting the electrical conductivity across the graphite slab, and assuming that the two outermost C-planes are connected to the two poles of a battery.
All simulations have been carried out using the Gromacsgromacs simulation package, version 2019.3. The analysis of trajectories has been made using home-made programs. These have been used to compute number and charge density profiles along the direction normal to the graphite planes, ion-ion radial distribution functions, self-diffusion coefficient based on the Einstein equation, orientation correlations among the ions and with respect to the planar IL/graphite interfaces. The electrostatic potential along the same normal direction has been computed from the charge density profile by integrating the electric field determined through Gauss theorem.
Figure 4: Snapshot of the ionic liquid part of Fig. 1 coarse grained through Eq. 1
A crucial aspect of the trajectory analysis is represented by the definition of dipoles within the ionic liquid, and their localisation in space. The first step in this process consists in identifying the position of each ion with its centre of charge:
${\bf R^{(Q)}_{\alpha}}=\frac{\sum_{i\in\alpha}\mid q_{i}\mid{\bf r_{i}}}{\sum_{i\in\alpha}\mid q_{i}\mid}$ (1)
whose definition corresponds to the centre of mass of a system of charges, whose mass is equal to the absolute value of their charge. This definition is not as generally accepted nor extensively used as that of the centre of mass. In this context, however, it provides a simple way to coarse grain the structure of [bmim][NTf2], which is fairly complex and difficult to visualise and interpret in its atomistic detail. For instance, the configuration of cations and anions corresponding to the snapshot in Fig. 1 is shown in Fig. 4, highlighting the drastic simplification of the IL structure achieved by the coarse graining expressed by Eq. 1
The method to highlight subtle correlations in the distribution of ions in the simulated trajectories is somewhat complex and consists of a few steps. For this reason, the partition of ions into neutral pairs is defined in the following Sec. III using actual simulation results to illustrate this protocol.
III Results
III.1 Properties of the homogeneous samples
Homogeneous samples of dry and water-contaminated [bmim][NTf2] have been simulated to validate the force field and to estimate baseline values for density, diffusion constant of cations and anions, and basic structural properties such as the radial distribution functions of cations, anions and water. Each sample consisted of $3072$ neutral ([bmim]+ and [NTf2]-, see Fig. 3) ion pairs, and up to $640$ water molecules in a cube periodically replicated in space.
Since the experimental density was fitted by increasing the LJ radii, the model reproduces the experimental density. Its temperature dependence is slightly underestimated. At the $20$ molar % concentration, water appears to be soluble in the re-parametrised [bmim][NTf2] model, in agreement with experiments.
Figure 5: Number (blue line) and charge (red line) density profiles across the uncharged ($\sigma=0$) capacitor.
The results show that the self-diffusion coefficient of cations and anions in pure, 3D homogeneous [bmim][NTf2] at $T=300$ K, although higher than that of the original Gromos 54a7 model, is still significantly below the experimental value measured by NMR.perk Moreover, and possibly more importantly, the estimated value depends on the time scale of the simulation, revealing a slightly sub-diffusive behaviour. Nevertheless, the mean square displacement is largely linear over a >50>50 ns interval, with an estimated value of $D$ of the order of $10^{-9}$ cm2/s for both anions and cations to within a large statistical error of $30$ %. This computed value has to be compared to the experimentally measured $10^{-7}$ $10^{-9}$ cm2/s, see the diffusion constant of [bmim][NTf2] measured by NMR as reported in Ref. perk, . We accept this discrepancy since unpolarizable ions are still the only way to simulate system sizes in excess of $10^{5}$ atoms at an acceptable cost. Addition of water at concentration as low as $1$ wgt% hardly affects the mobility of the ions. The self-diffusion coefficient of water in [bmim][NTf2] at 20 wgt% concentration and T=300T=300K turns out to be $(2.5\pm 0.5)\ 10^{-7}$ cm2/s, i.e., two orders of magnitude lower than in pure water at the same $T$, $P$ conditions.
Figure 6: Number (blue line) and charge (red line) density profiles across the capacitor with $\sigma=2$ e/nm2 surface charge density. Notice the change of scale with respect to Fig. 5.
III.2 Structural, dynamic and electrostatic properties of the graphite / [bmim][NTf2] interface
The primary focus of the simulations is a planar capacitor consisting of $3072$ ([bmim]+, [NTf2]-) neutral ion pairs, enclosed in between two electrodes that are represented by the opposite faces of a graphite slab, whose hexagonal planes are parallel to the interface. This basic sample is illustrated in Fig. 1. Each simulation cell is orthorhombic, and excluding the graphite slab, each IL sample is nearly cubic.
The reference frame has been oriented in such a way that the interface is orthogonal to the $x$ axis. For this reason, local properties such as density and electrostatic potential profiles are plotted as a function of $x$. To simulate charged capacitors, small charges of up to a few $10^{-2}$ e (where e is the H+ charge) have been equally distributed among all the atoms of the two most external layers of graphite, qualitatively reflecting the picture provided by ab-initio (density functional theory) computations, see charging, . The charges are of equal magnitude and opposite sign on the two charged layers. Since also the IL slab is globally neutral, the whole sample is neutral.
Figure 7: Electric field and electrostatic potential as a function of the $x$ coordinate across the capacitor for two different values of the surface charge $\sigma$.
The charge profile across the capacitor is the required ingredient to compute the electrostatic potential as a function of $x$, displayed in Fig. 7 for the $T=300$ K sample at $\sigma=0$ and at $\sigma=2$ e/nm2. As expected, the electrostatic potentials is flat and the electric field vanishes over a wide central region of the capacitor, independently from the surface charge.
The output of the electrostatic potential analysis will provide data to compute the differential capacitance per unit area of the interface, defined as:
$C_{d}=\left(\frac{\partial\sigma}{\partial\varphi}\right)_{T,P}$ (2)
that represent a crucial parameter characterizing the electrostatic double layer at the interface, measuring the width of the double layer itself.
III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
The results of the previous subsections show that the average electric field vanishes away from the interfaces. The presence of Stark effects sufficiently strong to be detected by spectroscopy, however, does not require a non-vanishing static electric field. A detectable Stark effect is in fact due to a change in the average of the square modulus $E^{2}$ of the electric field acting on each ion. Such a quantity does not vanish even in bulk [bmim][NTf2], and changes in its value can result from subtle correlations compatible with the strict screening sum rules that exclude net charge separation and non-vanishing average electric field away from the interface. Thus, excluding the violation of these rules, the present investigation focuses on possible correlations among electrostatic dipoles.
Such an analysis can follow a variety of leads, but two aspects need to be retained. First, one needs to define dipoles inside an ionic fluid, which is certainly a challenging undertaking since charges in [bmim][NTf2] are mobile. Second, each dipole needs to have a compact geometry in such a way that its position in space can be specified with good precision, allowing the meaningful computation of a local density of dipoles everywhere through the system as well as the definition of distance dependent correlation functions. The subdivision of the system charges into dipoles cannot but also does not need to be permanent, but it will evolve with the changing configuration of ions in the ionic liquid. However, to lend credibility to the physical relevance of these dipoles, their lifetime should be long. Moreover, the probability distribution for the modulus $R_{d}$ over the dipoles population should be relatively narrow, and correspond to a charge separation comparable to the average distance of ions as given by a ion-ion pair distribution function.
With these considerations in mind, we set out to partition the moving charges of the system into a population of dipoles. The dipole moment of a set of charges is unambiguously defined provided the aggregation of charge is overall neutral. Therefore, we focus our attention on the dipole of neutral ion pairs, consisting of the association of one cation and one anion. The difficulty with this approach is that the way to partition an assembly of ions into a collection of dipoles by joining a [bmim]+ cation and an [NTf2]- anion is not unique. The most intuitive approach could be to join each ion with its nearest neighbour of the opposite sign. However, the property of ion A+ being the nearest neighbour of B- is not symmetric, since the nearest cation of B- might be C+, different from A+. Simple algorithms like starting from a given ion, joining it to the closest counter-ion not already engaged into a dipole do not solve the problem. At first, this approach provides a series of tightly bound dipoles, satisfying the requirements listed above. Soon, however, going down the series of cations, the inherent frustration in the process of pairing neighbouring ions becomes manifest since the algorithm increasingly joins pairs separated by long distances, making meaningless the location of the dipole in space.
To correct this ambiguity, one has to define the cation-anion pairing in some physically motivated way, that makes the choice unique and close to the intuitive idea of the fluid made of nearly equivalent neutral ion pairs of relatively short separation, which are easily located in space.
The approach that is proposed and tested here is to define such a unique partition of the ion population into dipoles by requesting that their total self-energy is a minimum among all possible subdivision of all cations and anions into neutral pairs.
Figure 8: Snapshots of the dipole configurations at the two extreme temperatures of the simulated annealing. The $x$ axis is orthogonal to the two IL - graphite interfaces, with graphite not shown in the picture. Dark gray arrows point to the left, red arrows point to the right. $T$ is a fictitious temperature introduced to drive the MC optimisation of the ions’ pairing into dipoles, not to be confused with the physical temperature (T=300 K) of the graphite - liquid [bmim][NTf2] simulation. The units of the fictitious temperature are discussed in the text.
Let us first consider the case of neutral graphite surfaces. The case of charged interfaces will require some slight modifications. The self energy of a dipole made of charges (q,q)(q,-q) separated by a distance $s$ usually is given as $u_{s}=-q^{2}/s$. In the present case, in which cations and anions consists of several atoms joined into a complex ion, the self energy of the dipole made of cation $\alpha$ and anion $\beta$ is defined as:
$u_{s}(\alpha,\beta)=\sum_{i\in\alpha;j\in\beta}\frac{q_{i}q_{j}}{|{\bf r_{i}^{\alpha}}-{\bf r_{j}^{\beta}|}}$ (3)
where $i$ and $j$ label atoms (of charges $q_{i}$ and $q_{j}$, respectively) belonging to cation $\alpha$ and anion $\beta$, respectively. To carry out the unbiased subdivision of any given configuration of ions into a configuration of dipoles, a procedure has been devised which:
  • starts from the pairing of each cation with its nearest anion which is not already engaged in another pair;
  • minimize the total self-energy of the assembly of dipoles by a Monte Carlo procedures, in which neighbouring pairs of neutral ion pairs (i.e., two cations and two anions) interchange their bonding. In other terms, given the two pairs $(i_{1},j_{1})$ and $(i_{2},j_{2})$, (fictitious) dipole bonds are swapped into $(i_{1},j_{2})$ and $(i_{1},j_{2})$ according to the energy variation:
    $\delta e=u_{s}(i_{1},j_{2})+u_{s}(i_{2},j_{1})-u_{s}(i_{1},j_{1})-u_{s}(i_{2},j_{2})$ (4)
Acceptance - rejection of the bond interchange is decided by the usual Metropolis rule at fictitious temperature $T$, whose progressive reduction (simulated annealing) drives the system towards the minimum. The scale of he temperature interval and the schedule of the annealing have been decided by trial and error using a short sequence of preliminary computations. Measuring charges in units of $e$ and distances in nm (a unit dipole in these units is equal to 48 Debye units), temperature has been decreased from $T=2$ to $T=0.025$ energy units in 80 runs of $64\ 10^{6}$ MC steps, each consisting of the attempted interchange of bonds between neighbouring dipoles. Since the attempted MC move is a discrete event that cannot be fractioned, the acceptance ratio of the MC cannot be tuned, but it exceeded $0.5$ at high $T$, and decreased to $5\ 10^{-5}$ (one accepted move every 20 thousand moves) at the lowest $T=0.025$.
Figure 9: Dipoles’ self energy as a function of $T$ during the annealing.
The effect of the MC annealing on the definition of dipoles on top of a single configuration of the ions is illustrated in Fig. 8. Arrow represent the strength and spatial location of each dipole in the sample. The choice of the location of each dipole is explained in the following paragraph. The bottom panel shows the final configuration resulting from the simulated annealing. The configuration used as an example has been extracted from the simulation of the capacitor at $T=300$ K, $P=1$ bar, $\sigma=0$. The topmost panel shown a configuration of dipoles selected at random during the simulation at the highest temperature ($T=2$) of the simulated annealing protocol. The color of each arrow is assigned as follows: if the $x$ component of the dipole points to the right, the arrow is painted red, otherwise it is painted gray. The number of red dipoles equals the number of gray dipoles to within $1$ %. From the figure, it is apparent that at the end of the MC procedure, dipoles are of similar size and the sum of their self-energies, although unlikely to be the true absolute minimum, is already low in energy (see the self energy as a function of $T$ during the annealing in Fig. 9). It might be worth emphasising that this procedure is carried out at fixed position of all atoms in the system, only the distribution of fictitious dipole bonds among ions being interchanged.
Figure 10: Probability distribution for the charge separation of the dipoles at $T=0.025$ (end of the annealing) for the neutral interface (blue line) and for the charged interface at $\sigma=2$. The two distributions have been computed from the two dipole populations shown in Fig. 8. The area under the curve is normalised to: 100 in the case of the neutral interface; $100\times(3072+239)/3072$ for the charged interface, to account for the 239 ion pairs added to represent the charge on the electrodes.
The probability distribution for the separation $R_{d}$ of the charges for the dipoles in the final configuration is shown in Fig. 10. The average value $\langle R_{d}\rangle=0.5$ nm corresponds well to the geometric separation of cations and anions in the radial distribution function of the homogeneous [bmim][NTf2] liquid at the same conditions. The full width at half peak of about $0.1$ nm is sufficiently short to support the idea that the dipoles defined in this way are indeed physical associations of ions and not only fleeting geometrical details in the evolving ionic configuration.
Figure 11: Density of dipole as a function of position along the $x$ axis. The red lines show the $\langle d_{x}\rangle$ component, while the blue line shows the average $\langle d_{y}+d_{z}\rangle/2$ of the $d_{y}$ and $d_{z}$ components. The black rectangles represent the graphite electrodes. When $\sigma\neq 0$, the average of $\langle d_{x}\rangle$ does not vanish far from the interfaces.
The compact geometry of the ion pair carrying the dipole (say, dipole $\alpha$) allows us to define its geometric location through the relation:
$r^{dipole}_{\alpha}=\frac{\sum_{i\in\alpha}\mid q_{i}\mid{\bf r_{i}}}{\sum_{i\in\alpha}\mid q_{i}\mid}$ (5)
This allows to compute the density of dipole moment along each direction, which is displayed in Fig. 11 (a) for the sample illustrated in Fig. 8, having $\sigma=0$ surface charge density on the graphite electrodes. It is apparent that the distribution of dipole moments vanishes over most of the sample, with possibly a narrow peak in proximity of the two interfaces.
It can be supposed that a variety of subdivisions into pairs have comparable and relatively low self-energy, all representing the properties of the ideal unique and best pairing choice. However, repeated pairing optimisations, carried out by starting from different randomized starting point had a minimum fraction of equal pairs at the end of the simulated annealing of 66 %, showing that the self-energy basin around the (unknown) ideal minimum is relatively wide, but also well defined, without a plurality of significantly different local minima making ambiguous the pairing of ions into neutral pairs.
The visual inspection of the dipole configuration in Fig. 8 does not show any clear regularity. A more quantitative analysis involving the determination of the structure factor for the mixture of red and gray dipoles might still reveal the presence of clusters of dipoles pointing towards the same interface, providing at the same time an estimate of the size of these hypothetical domains.
The same protocol could be used to pairs ions into dipoles when the capacitor plates carry a non vanishing surface charge density. The resulting configuration of dipoles, however, is less regular than the one obtained at $\sigma=0$, and the MC optimisation is less able to provide a low energy partition of ions in dipoles of nearly homogeneous length. The reason is that the excess of ions of a given $\pm e$ charge screening the $\mp\sigma$ charge on the plate cannot find a neibouring partner in their immediate vicinity, being forced to form dipoles with ions located far into the IL bulk, or even close to the opposite plate, there an excess of counter ions can be found. In fact, at the end of virtually every simulated annealing cycles, the dipoles population included a few long separation (high self-energy) dipoles, unable to escape a local minimum of the total self-energy, and only a sustained effort is able to cure these configurations through very long simulated annealing cycles.
To overcome this problem, the original ions’ population is supplemented by further ions, whose role is to represent the charge on the capacitor plates, providing a pairing opportunity for the counter-ions that populate the first IL layer next to the interface interface. This procedure is better illustrated with an example. Let us consider the same IL sample of $3072$ ion pairs, enclosed in between two graphite electrodes whose interfacial layers (each consisting of $4784$ carbon atoms) now carry $\pm 0.05\ e$ for a total of $\pm 239.2$ e on the two capacitor plates. In the simulation, the left plate is negatively charged, and the rightmost plate is positively charged. Since each carbon atoms carries only $\pm 0.05$ e, the representation of the charge on each electrode cannot be obtained by including all these atoms in the dipole formation protocol. Instead, a number of integer charges corresponding to the total charge on each electrode is introduced. In the present example $239$ e-e charges are located on the leftmost electrode, and 239 +e+e charges are positioned on the rightmost electrode.
At the beginning, cations are sorted in order of increasing distance from the anode. Then, each of the $239$ cations closest to the anode, of coordinates $\{{\bf r_{i}},i=1,...,239\}$ is paired with a e-e charge located $0.2$ nm below the graphite plane, at the same $(y_{i},z_{i})$ in-plane position of the corresponding cation. The specular operation is carried out with the opposite polarity electrode, interchanging the role of cations and anions, using the distance from the positively charged plate to select the anions to be paired with the surface positive charge.
After this first $239+239$ pairs are formed, linking ions of opposite sign across the interface, the remaining 2594 ion pairs are given an initial pairing into dipoles precisely as in the $\sigma=0$ case. Then, the MC annealing is carried out following the same schedule as in the neutral case. Bond interchanges concern both the original ions, and the ions added to represent the surface charge.
Following this approach, the optimization of the self energy runs much smoothly and as effectively as in the neutral case, producing a probability distribution of dipole separations $R_{d}$ virtually superimposable to the one from the neutral case. The only difference is a tail at short separations apparent in the $\sigma=2$ e/nm2 plot, due to the tight pairs joining the added surface charges to the counter-ions in the inner screening layer (see Fig. 10). In the annealed configuration, 85 % of the dipoles point towards the right electrode. As apparent from the figure, the dipole density has a net excess of $d_{x}$ dipole moment along the entire capacitor along $x$. The net sum of all dipoles along $x$ has a direction which is such to partly balance the dipole formed by the two graphite surfaces of $\pm\sigma$ charge density. The observation of an excess of dipole moment far from charged interfaces is the major result of the present investigation.
Also in this case, repeating the simulated annealing after randomising the pairing of ions results in a set of ion pairs whose overlap with any previous one is very sizeble. In fact, the overlap seems to be even stronger than in the neutral case, perhaps suggesting that the minimum energy basin is slightly narrower and somewhat simpler for charged interfaces than for the neutral ones.
Figure 12: Dipoles’ configuration at he end of the simulated annealing optimization as a function of $\sigma$. Red dipoles point to the right side of the figure, white dipoles point to the left. The prevailing direction of dipoles in the charged interface case is such to partially compensate the dipole moment of the two charged graphite layers forming the plates of the capacitor.
A quantitative assessment of the dipoles’ alignemenmt along $x$ is given in Fig. 11, showing the average dipole moment along the directions parallel andnormal to the interface. The data reported in this figure refer to which is displayed in Fig. 11 (b) for the sample illustrated in Fig. 12, having $\sigma=2$ surface charge density on the graphite electrodes. It is apparent that the distribution of the $d_{x}$ component of the dipole moment does not vanish over the entire with of the capacitor.
IV Summary and discussion
The interface between ionic liquids and solid electrodes is a fascinating subject that joins the current interest for a vast number of electrochemical applications with concepts and theories dating back to one century ago.korn1 ; korn2 ; korn3 ; bock . This fertile combination, together with the complex structure of ions and the nanostructures texture of ionic liquids, has fostered a variety of results of practical interest, and also some surprises that challenge the accepted picture of the electrified interfaces. Among the aspects that are being challenged is the long held notion that, beyond a narrow interfacial region of microscopic width, a Coulombic fluid in contact with a charged interface is virtually bulk-like. A strong indication in this direction is provided by the recent observationexpLangmuir of the Pockel effect at the interface between the [bmim][NTf2] ionic liquid and two planar solid parallel electrodes connected to a generator mantaining an electrostatic potential energy difference between the two sides of the IL slab. Upon charging the interface, a shift in the Raman frequencies was observed, arising from all Raman-active modes of cations and anions, independently from their distance from the interface.
To investigate this phenomenon, the interface between the same [bmim][NTf2] ionic liquid and a graphite electrode has been simulated by molecular dynamics, using an atomistic semi-empirical force field. An innovative analysis of trajectories allows to join [bmim]+ and [NTf2]- ions into neutral pairs, whose dipoles show an apparent alignement with the external electric field driven by the charge deposited on the graphite surface. This alignment is observed even well inside the IL bulk-like region, where the external field is exactly cancelled (on average) by the screening at the interface by the mobile ions in the IL. Since dipoles do not obey the strict screening conditions of charges, the observation does not violate any accepted theory of electrified interfaces, but opens the way to a whole new class of medium-long range correlations that remain hidden when one considers the system only as a collection of free ions.
Up to now, this observation concerns a geometric property in the distribution of ions throughout the system, that leaves open questions concerning its origin and its significance with respect to the observed effect on the Raman spectra measured on the experimental samples. The main question on its origin concerns the role of the shape anisotropy of the ions in transferring the information on the state of the interface deep into the IL bulk. The main question on the relevance of the geometric property concerns whether this correlation among virtual dipoles can alter the vibrational properties of the ionic liquid. The mechanism joining the geometric correlation with the change of Raman frequences has to respect the fact that the average electric field well inside the ionic liquid vanishes, independently from the charge on the electrode. This fact, however, does not prevent changes in the average square amplitude of the fluctuating electric field inside the ionic liquid, that could indeed give origin to the observed change of vibratioal frequences.
Because of these open questions that are currently being investigated, the present document represents a report on work in progress, that will be followed by a full manuscriprit on the comprehensive investigation of the [bmim][NTf2] / graphite interface and on the hidden correlations in the distribution of ions arising upon chaarging the electrodes.